OpenStudy (anonymous):

solve for x in the interval [0,2pi] 2sinx^2+cosx=1

OpenStudy (turingtest):

use the identity\[\sin^2x=1-\cos^2x\] then the cosines can be set to zero and factored...

OpenStudy (turingtest):

\[2\sin^2x+\cos x=1\]identity:\[\sin^2x=1-\cos^2x\] sub in\[2(1-\cos^2x)+\cos x=1\to 2\cos^2x-\cos x-1=0\]\[(2\cos x+1)(\cos x-1)=0\]\[2\cos x+1=0\to\cos x=-\frac{1}{2}\to x=\frac{2\pi}{3},\frac{4\pi}{3}\]\[\cos x-1=0\to\cos x=1\to x=0,2\pi\]so\[x=\left\{ 0,\frac{2\pi}{3},\frac{4\pi}{3},2\pi \right\}\]

OpenStudy (ash2326):

(sin x)^2=1-(cos x)^2 put this in the equation 2-2(cos x)^2+cos x= 1 let cos x =y 2-2y^2+y=1 2y^2-y-1=0 find factor of -2 such that their difference is -1 -2 and 1 2y^2-2y+y-1 2y(y-1)+1(y-1)=0 (2y+1)(y-1)=0 y=-1/2 or y=1 cos x= -1/2 or 1 x= 0, 2pi cos x=-1/2 x = 2pi/3 or 4pi/3

OpenStudy (anonymous):

i got soo close! thanks guys